# Resonances and resonance expansions for $\delta$-interactions on the half-space

### v2

@@ 2. Preliminaries: Point Interaction on the half-space
From (2.1) and (2.2) we see that the resolvent operators of point interactions on the half-space are rank one perturbations of the unperturbed resolved operators. So the point spectrum of $-\Delta_{\alpha,y}^{\textup{D}}$ and $-\Delta_{\alpha,y}^{\textup{N}}$ consists of at most one point. The discrete spectrum of point interactions on domains is studied in a more general setting in the forthcoming paper of the same authors [38].
@@
One should notice that as a comparison of the structure of the resolvents shows, it is not possible to obtain the operators $-\Delta_{\alpha,y}^{\textup{D}}$ and $-\Delta_{\alpha,y}^{\textup{N}}$ combining several point interactions in $\mathbb{R}^{3}$ and suitably choosing their distances and parameters $\alpha$ (see [2] Chapter II.1, Theorem 1.1.4).
3. Resonances of point interactions in the half-space
A first result of this work concerns the localization of the resonances for the operators $-\Delta_{\alpha,y}^{\textup{D}}$ and $-\Delta_{\alpha,y}^{\textup{N}}$. Let now $H$ be either $-\Delta_{\alpha,y}^{\textup{D}}$ or $-\Delta_{\alpha,y}^{\textup{N}}$. The resolvent operator of $H$ is defined as a bounded operator for $\operatorname{Im}z>0$. From the explicit expression given in (2.1) and (2.2), we see that it can be continued meromorphically to $\mathbb{C}$ This meromorphic continuation is completely determined by the corresponding resolvent kernel.
@@ 3. Resonances of point interactions in the half-space
A resonance for $H$ is a pole of the resolvent kernel of $(H-z^{2})^{-1}$ in $\left\{z\in\mathbb{C}\ \text{s.t}\ \operatorname{Im}{z}<0\right\}\cup\left\{0\right\}$.
@@
It is not excluded that $z=0$ could be a resonance.
@@
Due to the condition $\operatorname{Im}z<0$ a non vanishing resonance belong to the second sheet of the Riemann surface of the meromorphic operator function $(H-z^{2})^{-1}$ or of the corresponding integral kernel.
@@
We recall that the Dirichlet Laplacian in the half-space has no resonances at all, while the Neumann laplacian on the half-space has a zero energy resonance (the constant function). On the other hand the point interactions $\Delta_{\alpha,y}$ on the whole $\mathbb{R}^{3}$ have, whatever be $y$, at most a single resonance, existing only if $\alpha\geq 0$ and given in that case by $z=-4\pi i\alpha$.
Remark 3.4.
Distributional solutions of the equation $(H-z^{2})\psi_{z}=0$ are called resonance functions. Resonance functions belong locally to the operator domain, and for $z\neq 0$ they are exponentially growing at infinity. For $z=0$ two possibilities arise. The first is that the resonance function $\psi_{0}$ is actually an eigenvector, so belonging to the operator domain and corresponding to a zero resonance which is also an eigenvalue. The second is that $\psi_{0}\notin L^{2}$, it is a genuine resonance function and in this case behaves at infinity as the fundamental solution of the laplacian: $\psi_{0}\sim\frac{c}{|x-y|}$ (plus possibly a constant).
The properties of resonances of the point interaction $-\Delta_{\alpha,y}^{\textup{D}}$ are completely and explicitly described by the following Proposition.
@@
Consider the operator $-\Delta_{\alpha,y}^{\textup{D}}$ Then the following properties hold:
@@
    For all $\alpha\in\mathbb{R}$ and $y\in\mathbb{R}_{+}^{3}$, in every region of the form $\frac{k\pi}{y_{3}}<\left\lvert\operatorname{Re}{z}\right\rvert<\frac{(2k+1)\pi}{2y_{3}}$, with $k=1,2,\dots$, there are exactly two values of $z$, with opposite real part, such that $z^{2}$ is a resonance for $-\Delta_{\alpha,y}^{\textup{D}}$.
@@
    (3.1) $\operatorname{Im}{z}=\frac{1}{2y_{3}}\ln{\Biggl{(}\frac{\sin{(2y_{3}\operatorname{Re}{z})}}{2y_{3}\operatorname{Re}{z}}\Biggr{)}}.$
@@
    The value $z=0$ corresponds to a resonance if and only if $\alpha=-\frac{1}{8\pi y_{3}}$ and in that case it is an eigenvalue.
@@
    There is a $z$ on the negative imaginary semi-axis generating a resonance if and only if $\alpha<-\frac{1}{8\pi y_{3}}$.
@@
    In the region $\operatorname{Re}{z}\in\Bigl{(}-\frac{\pi}{2y_{3}},\frac{\pi}{2y_{3}}\Bigr{)}$ there are exactly two values $z$ which generate resonances if and only if $\alpha>-\frac{1}{8\pi y_{3}}$. They have opposite real part.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl{(}\frac{\pi}{2}+k\pi\Bigr{)}}$ for some non-negative even $k$, then $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl{(}\frac{\pi}{2}+k\pi\Bigr{)}}$ generates a resonance.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl{(}-\frac{\pi}{2}-k\pi\Bigr{)}}$ for some negative odd $k$, then $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl{(}-\frac{\pi}{2}-k\pi\Bigr{)}}$ generates a resonance.
@@
Consider the operator $-\Delta_{\alpha,y}^{\textup{N}}$. Then the following properties hold:
@@
    For all $\alpha\in\mathbb{R}$ and $y\in\mathbb{R}_{+}^{3}$, in every region of the form $\frac{(2k+1)\pi}{2y_{3}}<\left\lvert\operatorname{Re}{z}\right\rvert<\frac{(k+1)\pi}{y_{3}}$ with $k=0,1,\dots$, there are exactly two values $z$, with opposite real part, such that $z^{2}$ is a resonance for $-\Delta_{\alpha,y}^{\textup{N}}$. Moreover
    (3.2) $\operatorname{Im}{z}=\frac{1}{2y_{3}}\ln{\Biggl{(}-\frac{\sin{(2y_{3}\operatorname{Re}{z})}}{2y_{3}\operatorname{Re}{z}}\Biggr{)}}.$
@@
    The value $z=0$ corresponds to a resonance if and only if $\alpha=\frac{1}{8\pi y_{3}}$ and it is not an eigenvalue.
@@
    There is a $z$ on the negative imaginary semi-axis generating a resonance if and only if $\alpha>\frac{1}{8\pi y_{3}}$.
    When it exists, this $z$ is unique.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl{(}\frac{\pi}{2}+k\pi\Bigr{)}}$ for some positive odd $k$, then $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl{(}\frac{\pi}{2}+k\pi\Bigr{)}}$ generates an additional resonance.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl{(}-\frac{\pi}{2}-k\pi\Bigr{)}}$ for some negative even $k$, then $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl{(}-\frac{\pi}{2}-k\pi\Bigr{)}}$ generates an additional resonance.
@@
The resonances on the negative imaginary axis, existing under the conditions stated in Proposition 3.1(ii) and 3.2(ii) are usually called, especially in the physical literature, antibound states.
In the following we prove Proposition 3.1. The proofs of Proposition 3.2 is not reported because, barring some minor details, it follow the same path of the Dirichlet case.
A plot of Dirichlet resonance with $\operatorname{Re}z>0$ is given in Figure 1.
@@
We only give the proof for $iii)$, the proof of $i)$ and $ii)$ being entirely elementary. The derivative of $h$ is
@@
We claim that $h^{\prime}(a)\geq 0$ $\forall a\geq 0$ (where defined). This is a consequence of the elementary inequality
@@
being satisfied $\forall t\geq 0$, with the equality only valid in $t=0$. The continuity of $h$ and its derivative in every interval of the form $\Bigl{(}\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr{)}$ with $k=0,1,2,\dots$, then implies that $h$ is monotone increasing there. The second derivative $h^{\prime\prime}$ is given by
@@
It holds that $h^{\prime\prime}>0$ when $t^{2}+\sin^{2}{t}\cos^{2}{t}+2\sin^{2}{t}-4t\sin{t}\cos{t}>0,$ with $t=2y_{3}a$. This expression is greater or equal than $t^{2}-2t$, which is positive for $t>2$. So $h^{\prime\prime}(a)>0$ $\forall a>1/y_{3}$ in its domain. Since $1/y_{3}<\pi/y_{3}\leq k\pi/y_{3}$ for $k=1,2,\dots$, then $h^{\prime\prime}>0$ in each interval $\Bigl{(}\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr{)}$ with $k=1,2,\dots$ and it is convex there. ∎
Proof of Proposition 3.1.
The definition of resonance and Eq. 2.1 imply that the non real resonances of $-\Delta_{\alpha,y}^{\textup{D}}$ are determined by the solutions of the equation $\Gamma_{\alpha,y}^{\textup{D}}(z)=0$ with $\operatorname{Im}{z}<0$. Hence, we set $z=a+ib$ in $\Gamma_{\alpha,y}^{\textup{D}}(z)=0$, with the prescription $b<0$. This way we get
@@
(3.3) $\begin{cases}-\alpha-\frac{b}{4\pi}=e^{-2y_{3}b}\frac{\cos{(2y_{3}a)}}{8\pi y_{3}}\\ \frac{a}{4\pi}=e^{-2y_{3}b}\frac{\sin{(2y_{3}a)}}{8\pi y_{3}}\end{cases}.$
We start by looking for solutions on the negative imaginary semi-axis. So we set $a=0$. We observe that in this case the second equation is satisfied $\forall b$. The first equation of the system is then
@@
Since $\lim_{b\to-\infty}f(b)=+\infty$ and $f$ is a continuous monotonically decreasing function, a solution in $(-\infty,0]$ exists only if $f(0)<0$. This happens when $\alpha\leq-(8\pi y_{3})^{-1}$, otherwise there are no resonances on the negative imaginary semi-axis. The possible solution is in fact unique because $f^{\prime}(b)=(1-e^{-2y_{3}b})/(4\pi)<0$ $\forall b<0$. This proves part iii) of the Proposition. In particular, one has $b=0$ if and only if $\alpha=-(8\pi y_{3})^{-1}$. So, $z=0$ is a resonance when this condition is satisfied. In this case, it is easy to check that the resonance function
(3.4) $\psi_{0}(x)=\frac{1}{4\pi\left\lvert x-{y}\right\rvert}-\frac{1}{4\pi\left\lvert x-\overline{y}\right\rvert}\in L^{2}(\mathbb{R}^{3}_{+})$
and it is an eigenvector, and $z=0$ is an eigenvalue. This proves part ii).
Another case of interest in the study of Eq. 3.3 is when $a=\frac{k\pi}{2y_{3}}$, for some $k=\pm 1,\pm 2,\dots$. In this case it is easy to see that the system does not admit any solution, because the right side of the second equation of the system is zero, while the left one isn’t. So there are no resonances on the lines $\operatorname{Re}{z}=\frac{k\pi}{2y_{3}}$, $k=\pm 1,\pm 2,\dots$ outside the real axis.
Also, since we already considered the $a=0$ case, we can rewrite the second equation of the system Eq. 3.3 as
@@
Notice that, as expected, no solution can exist on the upper half complex plane, barring the ones on the real axis and on the positive imaginary semi-axis, this being a consequence of the self-adjointness of the operator. Since the left side is always smaller than 1, a solution is possible only if $b<0$. It is immediate to see that there are no solution when $\pi+2k\pi<2y_{3}\left\lvert a\right\rvert<2\pi+2k\pi$, $k=0,1,2,\dots$, because in that case the left side would be negative.
We now consider the cases when $a=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}$ for some $k\in\mathbb{Z}$. Eq. 3.3 is then

### v3

@@ 3. Resonances of point interactions in the half-space
The previous definition agrees with the treatment based on [3] in the case of one or more point interactions. Due to the structure of the resolvent operator, it is also consistent with the treatment of Sjöstrand and Zworski in [50] and [49] (see also the treatise [28] by Dyatlov and Zworski). Notice that from the point of view of spectral theory the relevant object is $z^{2}$ and not directly $z$. In particular, a solution $z$ of Eq.˜3.3 such that $\operatorname{Im}{z}\geq 0$ and the corresponding solution $\psi$ of $H\psi=z^{2}\psi$ is in $L^{2}(\mathbb{R}_{+}^{3})$ gives rise to an eigenvalue $z^{2}$ of $H$. Due to self-adjointness, such $z$ can only belong to the real axis or to the positive imaginary axis. Moreover, still due to self-adjointness, for $\operatorname{Im}z>0$ the only possible solutions to Eq.˜3.3 correspond to eigenvalues. A solution $z$ of Eq.˜3.3 such that $\operatorname{Im}{z}\leq 0$ and not corresponding to an eigenvalue, gives rise to a genuine continuation resonance of $H$. We prefer however for the sake of simplicity to retain the spectral parameter $z$ as the main object. According to Definition˜3.1 there is no distinction between eigenvalues and continuation resonances. In the following, however, we will always make clear when a solution of Eq.˜3.3 gives rise to an eigenvalue or to a continuation resonance.
@@
We recall that the Dirichlet and Neumann Laplacian in the half-space have no resonances at all. On the other hand the point interactions $\Delta_{\alpha,y}$ on the whole $\mathbb{R}^{3}$ have, whatever be $y$, a single resonance, existing for every $\alpha\in\mathbb{R}$ and given in that case by $z=-4\pi i\alpha$. For $\alpha<0$ the resonance corresponds actually to the eigenvalue $(4\pi\alpha)^{2}$, while for $\alpha\geq 0$ is a genuine continuation resonance.
@@
The multiplicity of a resonance $z$ of the operator $H$ is the algebraic multiplicity of the corresponding solution of Eq.˜3.3.
@@
In particular, the resonances (eigenvalues or continuation resonances) are always simple for the operators here treated apart for the Dirichlet case when $\alpha=-1/(8\pi y_{3})$, where the multiplicity is two. For a resonance $z$ to have multiplicity greater than one it must hold that
$\begin{cases}\Gamma_{\alpha,y}(z)=0\\ \frac{\partial\Gamma_{\alpha,y}}{\partial z}(z)=0\end{cases}$
or explicitly
$\begin{cases}\alpha-\frac{iz}{4\pi}\pm\frac{e^{2iy_{3}z}}{8\pi y_{3}}=0\\ \pm e^{2iy_{3}z}=1\end{cases}$
where the plus sign denotes the Dirichlet case and the minus sign the Neumann one.
The second equation of the Dirichlet system has solutions $z_{\textup{D}}=k\pi/y_{3}$, $k\in\mathbb{Z}$. Substituting it in the first equation leads to the condition
$\alpha-\frac{ik}{4y_{3}}+\frac{1}{8\pi y_{3}}=0,$
which is verified if and only if $k=0$ (hence $z=0$) and $\alpha=-1/(8\pi y_{3})$.
The second equation of the Neumann system has instead solutions $z_{\textup{N}}=\pi/(2y_{3})+k\pi/y_{3}$, $k\in\mathbb{Z}$. Substitution easily shows that for thew system to be solvable it needs to be k=1/2k=-1/2 and hence there cannot be resonances with multiplicity grater than one in the Neumann case. Finally, noticing that $\frac{\partial^{2}\Gamma_{\alpha,y}}{\partial z^{2}}(z)\neq 0$ for all $z\in\mathbb{C}$ we conclude that the multiplicity of a resonance is at most two.
The spectrum of point interactions $\Delta_{\alpha,y}^{\textup{D}}$ and $\Delta_{\alpha,y}^{\textup{N}}$ is given in the following proposition. Other properties about the spectrum of point interaction in more general unbounded sub-domains of $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ will appear in the forthcoming article [45] by these authors.
Proposition 3.2.
The spectra of the operators $\Delta_{\alpha,y}^{\textup{D}}$ and $\Delta_{\alpha,y}^{\textup{N}}$ are characterized as follows.
  • a)
    $\sigma_{c}(-\Delta_{\alpha,y}^{\textup{D}})=\sigma_{ac}(-\Delta_{\alpha,y}^{\textup{D}})=[0,+\infty)$. Moreover, $\sigma_{p}(-\Delta_{\alpha,y}^{\textup{D}})$ is empty or it consists of a single non positive simple eigenvalue existing if and only if $\alpha<-\frac{1}{8\pi y_{3}}$.
  • b)
    $\sigma_{c}(-\Delta_{\alpha,y}^{\textup{N}})=\sigma_{ac}(-\Delta_{\alpha,y}^{\textup{N}})=[0,+\infty)$. Moreover, $\sigma_{p}(-\Delta_{\alpha,y}^{\textup{N}})$ is empty or it consists of a single simple negative eigenvalue existing if and only if $\alpha<\frac{1}{8\pi y_{3}}$.
Proof.
The general theory of self-adjoint extensions of positive symmetric operators with equal and finite deficiency indices assures that all the members of the family of self-adjoint extension share the same continuous spectrum, and that the number of negative eigenvalues of member of the family cannot exceed the deficiency index (counting multiplicity), which here is one (see for example [1], Section 85). The absence of singular continuous spectrum is a consequence of the explicit form of the resolvent given in formulae Proposition˜2.1 and Proposition˜2.2 and of the direct application of the standard criterion in [46], Theorem XIII.19. The quantitative threshold values of the existence of eigenvalues are obtained by inspection of formulae (3.8) with $a=0$ and $b\geq 0$ as given in the proof of the subsequent Proposition 3.4, to which we refer for a similar analysis. ∎
The properties of resonances of the point interactions $\Delta_{\alpha,y}^{\textup{D}}$ and $\Delta_{\alpha,y}^{\textup{N}}$ are completely and explicitly described by the following propositions.
Proposition 3.3.
Operators $\Delta_{\alpha,y}^{\textup{D}}$ and $\Delta_{\alpha,y}^{\textup{N}}$ have no continuation resonances on the real axis apart possibly $z=0$. Moreover, $z=0$ is a resonance if and only if $\alpha=-1/(8\pi y_{3})$ in the Dirichlet case and if and only if $\alpha=1/(8\pi y_{3})$ in the Neumann case.
Proof.
Let be $z\in\mathbb{C}$ such that $\Gamma_{\alpha,y}(z)=0$. Depending on the boundary condition, according to the form of $\Gamma_{\alpha,y}$ given in Proposition˜2.1 and Proposition˜2.2 the conditions takes the form (++ sign Dirichlet, - sign Neumann)
$\alpha-\frac{iz}{4\pi}\pm\frac{e^{2iy_{3}z}}{8\pi y_{3}}=0.$
To investigate the existence of real resonances we substitute $z=a\in\mathbb{R}$ in the above equation. Separating the real and imaginary part leads to the following system
$\begin{cases}\cos{(2y_{3}a)}=\mp 8\pi\alpha y_{3}\\ \sin{(2y_{3}a)}=\pm 2ay_{3}\end{cases}.$
Since $y_{3}>0$ the second equation admits only $a=0$ as possible solution. In this case, because of the first equation, it must be $\alpha=\mp 1/(8\pi y_{3})$. ∎
@@
Distributional solutions (where distributions are meant in $\mathbb{R}^{3}_{+}$) of the equation $(H-z^{2})\psi_{z}=0$ where again $H$ denotes either $\Delta_{\alpha,y}^{\textup{D}}$ or $\Delta_{\alpha,y}^{\textup{N}}$, are called resonance functions corresponding to the resonance $z$. Resonance functions belong to the space $H^{2}_{\text{loc}}(\mathbb{R}^{3}_{+}\setminus\{y\})$, satisfy the boundary conditions on the boundary plane $x_{3}=0$ and at the singularity $y$ (see remark 2.3), and for $\operatorname{Im}z<0$ they are exponentially growing at infinity. For $z=0$ it can be shown that resonance functions belong to certain weighted Sobolev spaces depending on the dimension and in particular they decay at infinity in dimension three (see [48] for the case of several point interactions, which in this respect is completely analogous to the present case, and the seminal paper [38] for the case of Schrödinger operators). However, will not need these informations in the present analysis.
@@
If $z=0$ is a solution of Eq.˜3.3 two possibilities arise. The first is that the resonance function $\psi_{0}$ is actually an eigenvector, so belonging to the operator domain and corresponding to a zero resonance which is also an eigenvalue. The second is that $\psi_{0}\notin L^{2}$, it is a genuine resonance function and in this case $\psi_{0}$ behaves at infinity as the fundamental solution of the Laplacian: $\psi_{0}\sim\frac{c}{|x-y|}$. These facts are explicit in subsequent formula (3.5) for the Dirichlet case where a threshold eigenvalue appears, and in the formula (3.7) for the Neumann case where a genuine threshold resonance appears.
@@
Consider the operator $\Delta_{\alpha,y}^{\textup{D}}$ . Then the following properties hold:
@@
    For all $\alpha\in\mathbb{R}$ and $y\in\mathbb{R}_{+}^{3}$, in every region of the form $\frac{k\pi}{y_{3}}<\left\lvert\operatorname{Re}{z}\right\rvert<\frac{(2k+1)\pi}{2y_{3}}$, with $k=1,2,\dots$, there are exactly two resonances $z$ for $\Delta_{\alpha,y}^{\textup{D}}$ with opposite real part.
@@
    $\operatorname{Im}{z}=\frac{1}{2y_{3}}\ln{\Biggl(\frac{\sin{(2y_{3}\operatorname{Re}{z})}}{2y_{3}\operatorname{Re}{z}}\Biggr)}.$ (3.4)
@@
    The value $z=0$ is a resonance if and only if $\alpha=-\frac{1}{8\pi y_{3}}$ and in that case it is an eigenvalue. The corresponding eigenvector is
    $\psi_{0}(x)=\frac{1}{4\pi\left\lvert x-{y}\right\rvert}-\frac{1}{4\pi\left\lvert x-\overline{y}\right\rvert}$ (3.5)
@@
    There is a resonance $z$ on the negative imaginary semi-axis if and only if $\alpha<-\frac{1}{8\pi y_{3}}$.
@@
    The region $\operatorname{Re}{z}\in\Bigl(-\frac{\pi}{2y_{3}},\frac{\pi}{2y_{3}}\Bigr)$ contains resonances if and only if $\alpha>-\frac{1}{8\pi y_{3}}$. In such a case there are two resonances with opposite real part.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl(\frac{\pi}{2}+k\pi\Bigr)}$ for some non-negative even $k$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl(\frac{\pi}{2}+k\pi\Bigr)}$.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl(-\frac{\pi}{2}-k\pi\Bigr)}$ for some negative odd $k$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl(-\frac{\pi}{2}-k\pi\Bigr)}$.
@@
Consider the operator $\Delta_{\alpha,y}^{\textup{N}}$. Then the following properties hold:
@@
    For all $\alpha\in\mathbb{R}$ and $y\in\mathbb{R}_{+}^{3}$, in every region of the form $\frac{(2k+1)\pi}{2y_{3}}<\left\lvert\operatorname{Re}{z}\right\rvert<\frac{(k+1)\pi}{y_{3}}$ with $k=0,1,\dots$, there are exactly two resonances $z$ for $\Delta_{\alpha,y}^{\textup{N}}$ with opposite real part. Moreover
    $\operatorname{Im}{z}=\frac{1}{2y_{3}}\ln{\Biggl(-\frac{\sin{(2y_{3}\operatorname{Re}{z})}}{2y_{3}\operatorname{Re}{z}}\Biggr)}.$ (3.6)
@@
    The value $z=0$ is a resonance if and only if $\alpha=\frac{1}{8\pi y_{3}}$ and in that case it is not an eigenvalue. The corresponding resonance function is
    $\psi_{0}(x)=\frac{1}{4\pi\left\lvert x-{y}\right\rvert}+\frac{1}{4\pi\left\lvert x-\overline{y}\right\rvert}$ (3.7)
@@
    There is a resonance $z$ on the negative imaginary semi-axis if and only if $\alpha>\frac{1}{8\pi y_{3}}$.
    When it exists this $z$ is unique.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl(\frac{\pi}{2}+k\pi\Bigr)}$ for some positive odd $k$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl(\frac{\pi}{2}+k\pi\Bigr)}$.
@@
    If $\alpha=\frac{1}{8\pi y_{3}}\ln{\Bigl(-\frac{\pi}{2}-k\pi\Bigr)}$ for some negative even $k$, then there is a resonance in $z=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}-\frac{i}{2y_{3}}\ln{\Bigl(-\frac{\pi}{2}-k\pi\Bigr)}$.
@@
The resonances on the negative imaginary axis, existing under the conditions stated in Proposition 3.4(iii) and 3.5(iii) are usually called, especially in the physical literature, antibound states.
In the following we prove Proposition˜3.4. The proof of Proposition˜3.5 is not reported because, barring some minor details, it follow the same path of the Dirichlet case.
A plot of Dirichlet resonances with $\operatorname{Re}z>0$ is given in Figure 1.
@@
We only give the proof for $iii)$, the proof of $i)$ and ii)ii) being entirely elementary. The derivative of $h$ is
@@
We claim that $h^{\prime}(a)\geq 0$ for all $a\geq 0$ (where defined). This is a consequence of the elementary inequality
@@
being satisfied for all $t\geq 0$, with the equality only valid in $t=0$. The continuity of $h$ and its derivative in every interval of the form $\Bigl(\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr)$ with $k=0,1,2,\dots$, then implies that $h$ is monotone increasing there. The second derivative $h^{\prime\prime}$ is given by
@@
It holds that $h^{\prime\prime}>0$ when $t^{2}+\sin^{2}{t}\cos^{2}{t}+2\sin^{2}{t}-4t\sin{t}\cos{t}>0,$ with $t=2y_{3}a$. This expression is greater or equal than $t^{2}-2t$, which is positive for $t>2$. So $h^{\prime\prime}(a)>0$ for all $a>1/y_{3}$ in its domain. Since $1/y_{3}<\pi/y_{3}\leq k\pi/y_{3}$ for $k=1,2,\dots$, then $h^{\prime\prime}>0$ in each interval $\Bigl(\frac{k\pi}{y_{3}},\frac{(2k+1)\pi}{2y_{3}}\Bigr)$ with $k=1,2,\dots$ and it is convex there. ∎
Proof of Proposition˜3.4.
The definition of resonance and Proposition˜2.1 imply that the continuation resonances of $\Delta_{\alpha,y}^{\textup{D}}$ are determined by the solutions of the equation $\Gamma_{\alpha,y}^{\textup{D}}(z)=0$ with $\operatorname{Im}{z}<0$. Hence, we set $z=a+ib$ in $\Gamma_{\alpha,y}^{\textup{D}}(z)=0$, with the prescription $b<0$. This way we get
@@
$\begin{cases}-\alpha-\frac{b}{4\pi}=e^{-2y_{3}b}\frac{\cos{(2y_{3}a)}}{8\pi y_{3}}\\ \frac{a}{4\pi}=e^{-2y_{3}b}\frac{\sin{(2y_{3}a)}}{8\pi y_{3}}\end{cases}.$ (3.8)
We start by looking for solutions on the negative imaginary semi-axis and we set $a=0$. We observe that in this case the second equation is satisfied for all $b$. The first equation of the system is then
@@
Since $\lim_{b\to-\infty}f(b)=+\infty$ and $f$ is a continuous monotonically decreasing function, a solution in $(-\infty,0]$ exists only if $f(0)<0$. This happens when $\alpha\leq-(8\pi y_{3})^{-1}$, otherwise there are no resonances on the negative imaginary semi-axis. The possible solution is in fact unique because $f^{\prime}(b)=(1-e^{-2y_{3}b})/(4\pi)<0$ for all $b<0$. This proves part iii) of the Proposition. In particular, one has $b=0$ if and only if $\alpha=-(8\pi y_{3})^{-1}$. So, $z=0$ is a resonance when this condition is satisfied. In this case, it is easy to check that the resonance function
$\psi_{0}(x)=\frac{1}{4\pi\left\lvert x-{y}\right\rvert}-\frac{1}{4\pi\left\lvert x-\overline{y}\right\rvert}$ (3.9)
belongs to $L^{2}(\mathbb{R}^{3}_{+})$; so it is an eigenvector and $z=0$ is an eigenvalue. This proves part ii).
Another case of interest in the study of Eq.˜3.8 is when $a=\frac{k\pi}{2y_{3}}$, for some $k=\pm 1,\pm 2,\dots$. In this case it is easy to see that the system does not admit any solution, because the right side of the second equation of the system is zero, while the left one isn’t. So there are no resonances on the lines $\operatorname{Re}{z}=\frac{k\pi}{2y_{3}}$, $k=\pm 1,\pm 2,\dots$ outside the real axis.
Also, since we already considered the $a=0$ case, we can rewrite the second equation of the system Eq.˜3.8 as
@@
Notice that, as expected, because of the self-adjointness of the operator, the solutions with $\operatorname{Im}{z}\geq 0$ can only be found on the real axis or on the positive imaginary semi-axis.
Since the left side is always smaller than 1, a solution is possible only if $b<0$. It is immediate to see that there are no solution when $\pi+2k\pi<2y_{3}\left\lvert a\right\rvert<2\pi+2k\pi$, $k=0,1,2,\dots$, because in that case the left side would be negative.
We now consider the cases when $a=\frac{\pi}{4y_{3}}+\frac{k\pi}{2y_{3}}$ for some $k\in\mathbb{Z}$. From formula Eq.˜3.8 we get